# On computational complexity and average-case hardness of shallow-depth Boson Sampling

### 05 Mar 2026 v3

@@ Appendix E About Cayley transform
where we denote $\bm{t}^{*}$ as an input configuration vector permuted by $\bm{P}^{-1}$ from the fixed input configuration $\bm{t}$ in Eq. (1), and $S_{N}$ is NN-mode permutation group. One can easily check that the common denominator for $\prod_{j=1}^{N}\left[V(\theta)_{\bm{s},\bm{t}^{*}}\right]_{\sigma_{j},j}$ is $\prod_{i=1}^{m}q_{i}(\theta)]^{N}$, and it does not change with permutation $\sigma$. Let us define $Q(\theta)=[\prod_{i=1}^{m}|q_{i}(\theta)|^{2}]^{N}$, which is a degree $4mN$ polynomial in $\theta$. Then, $Q(\theta)$ serves as the common denominator for the output probability. Hence, the output probability can be represented as $p_{\bm{s}}(V(\theta))=\frac{P(\theta)}{Q(\theta)}$, with $P(\theta)$ also a degree $4mN$ polynomial in $\theta$. ∎
@@ Appendix G Proof of Lemma 3
Let $\bar{p}_{\bm{s}}(C)$ be the output probability distribution from the approximate sampler $\mathcal{S}$ with the given linear optical circuit $C$. Also, let $\mathcal{C}_{M,N}$ be the set of collision-free outcomes of Boson Sampling, for mode number $M$ and photon number NN. Then $\bar{p}_{\bm{s}}(C)$ satisfies
@@
in polynomial time in NN and $\alpha^{-1}$ is in $\rm{BPP}^{\rm{NP}^{\mathcal{S}}}$. Using $\operatorname*{\mathbb{E}}_{\bm{s}\sim\mathcal{G}_{M,N}}[\bar{p}_{\bm{s}}(C)]=\binom{M}{N}^{-1}\sum_{\bm{s}\in\mathcal{C}_{M,N}}\bar{p}_{\bm{s}}(C)\leq\binom{M}{N}^{-1}$,
@@ Appendix H On average-case hardness including collision outcomes
More specifically, let the mode number $M=\Theta(N^{\gamma})$ for any $\gamma\geq 1$. Let us define $\mathcal{Q}_{M,N}$ as the uniform distribution over all the possible outcomes of Boson Sampling with $M$ modes and NN photons, such that each outcome $\bm{s}\sim\mathcal{Q}_{M,N}$ is an $M$-dimensional output configuration vector $\bm{s}=(s_{1},\dots,s_{M})$ with the constraint $\sum_{i=1}^{M}s_{i}=N$. Also, we denote the number of $s_{i}=1$ in $\bm{s}$ as the number of singletons throughout this appendix. Here, let us assume that there exists $N_{0}=cN^{\lambda}$ for a constant $c$ and $0<\lambda<1$ such that it is likely that the number of singletons is more than $N_{0}$; we later provide the intuition to justify this assumption at the end of this appendix.
Hereafter, we denote each $(M_{0},N_{0})$ and $(M,N)$ as the system sizes (mode number & photon number) for the worst-case hardness (i.e., Theorem 3) and average-case hardness (we aim to prove), respectively. More specifically, given a worst-case circuit $C_{0}$ in $M_{0}=\Omega(N_{0})$ mode circuit architecture $\mathcal{B}\mathcal{B}^{*}$ and a fixed $M_{0}$-mode and $N_{0}$-photon collision-free outcome $\bm{s}_{0}$ depicted in Theorem 3, by our assumption, one can find a photon number NN that satisfies $N_{0}=cN^{\lambda}$ for a constant $c$ and $0<\lambda<1$, and mode number $M=\Theta(N^{\gamma})$ for $\gamma\geq 1$, such that most of the outcomes over $\mathcal{Q}_{M,N}$ have at least $N_{0}$ singletons. In the following, we show a possible approach for worst-to-average-case reduction for general outcomes, by describing how one can estimate $p_{\bm{s}_{0}}(C_{0})$ to within desired imprecision given oracle access to the well-estimated values $p_{\bm{s}}(U)$ for average-case outcomes $\bm{s}\sim\mathcal{Q}_{M,N}$ and $M$-mode average-case circuits over $U\sim\mathcal{H}_{\mathcal{A}}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$.
@@
We highlight the necessary ingredients that need to be addressed for the successful worst-to-average-case reduction by the above process. First, to make sure that we can also well-estimate $p_{\bm{s}_{1}}(U_{0})$ with high probability over $\bm{s}_{1}$ via the oracle access, one needs to justify that the distribution of $\bm{s}_{1}$ is sufficiently close to the distribution $\mathcal{Q}_{M/2,N-N_{0}}$. To help understand how the distribution of $\bm{s}_{1}$ can be close to $\mathcal{Q}_{M/2,N-N_{0}}$, consider the limiting case that $N_{0}=0$. Here, given that $M$ is much larger than NN (but still can be $M=\Theta(N)$), sampling $\bm{s}_{1}$ is equivalent to sampling from $\mathcal{Q}_{M,N}$ and post-selecting outcomes that have all the non-zero elements in the first $M/2$ mode, which yields the uniform distribution $\mathcal{Q}_{M/2,N}$. Therefore, given that $M$ is much larger than NN and $N_{0}$ is sufficiently small relative to NN, we expect that the distribution of $\bm{s}_{1}$ would not largely deviate from the distribution $\mathcal{Q}_{M/2,N-N_{0}}$.
@@
We now present a toy model that could offer an insight that, for a large portion of outcomes of Boson Sampling, a sufficient number of singletons would exist. In the following, we present the classical “balls-into-bins" problem where we throw NN balls into $M$ bins uniformly at random, which is indeed the classical analogue of Boson Sampling. We show that, given that $M=\Theta(N^{\gamma})$ with $\gamma\geq 1$, it is highly unlikely that the number of bins getting exactly one ball is less than $cN^{\lambda}$ for a constant $c$ and $0<\lambda<1$.
Let us first denote the random variable $X_{i}$ as the number of balls in $i$-th bin (such that $\sum_{i=1}^{M}X_{i}=N$), when we throw NN balls into $M$ bins uniformly at random. Here, our goal is to verify that over the experiments, it is highly unlikely to observe the outcomes that the number of $X_{i}=1$ is smaller than $cN^{\lambda}$. Since all the variables $X_{i}$ are correlated due to the constraint $\sum_{i=1}^{M}X_{i}=N$, instead, we consider i.i.d. random variable $Y_{i}$ each drawn from Poisson distribution $P(\frac{N}{M})$, such that $\operatorname*{\mathbb{E}}[\sum_{i=1}^{M}Y_{i}]=N$ and $\Pr[\sum_{i=1}^{M}Y_{i}=N]=e^{-N}\frac{N^{N}}{N!}=\Theta(\frac{1}{\sqrt{N}})$. The motivation for using Poissonian random variables is that the distribution of $(Y_{1},\dots,Y_{M})$ conditioned on the sum $\sum_{i=1}^{M}Y_{i}=N$ is equivalent to the distribution of the balls-into-bins problem $(X_{1},\dots,X_{M})$ (see, e.g., Ref. [ross2014introduction]). Accordingly, if it is exponentially unlikely that the number of $Y_{i}=1$ is smaller than $cN^{\lambda}$ as we sample $(Y_{1},\dots,Y_{M})$, this indicates that it is still exponentially unlikely that the number of $X_{i}=1$ is smaller than $cN^{\lambda}$ as we sample $(X_{1},\dots,X_{M})$, given that the post-selection probability for $\sum_{i=1}^{M}Y_{i}=N$ is at most inverse polynomial (i.e., $\Theta(\frac{1}{\sqrt{N}})$).
@@
We conclude this appendix by introducing an alternative approach, based on the closely related result in [bouland2023complexity] that establishes a hardness argument for general Boson Sampling in the saturated regime. Note that our convention on singletons closely parallels the notion of “clicks" introduced in [bouland2023complexity], which refers to the number of modes that are occupied by a non-zero number of photons. Analogous to our analysis, their key observation is that, even in the saturated regime, most outcomes have a sufficiently large number of clicks. Based on this observation, they show that, given an NN-photon outcome $\bm{s}$ with collisions that has a sufficient number of clicks $N_{0}$, there exists a poly-time constructible matrix $A\in\{0,1\}^{N_{0}\times N}$ such that computing $\text{Per}(A_{\bm{s}})$ is worst-case #P-hard, for $A_{\bm{s}}$ being a submatrix of AA by repeating rows according to $\bm{s}$. Then they establish a worst-to-average-case reduction by perturbing AA into their average-case matrix and performing polynomial interpolation (for details, see [bouland2023complexity]).
Accordingly, since we employ a conceptually similar strategy in the main text, establishing worst-case hardness for shallow-depth circuits with a large enough number of clicks is sufficient to obtain an analogous hardness result for shallow-depth Boson Sampling. This requires constructing a shallow-depth circuit (as done in the proof of Theorem 3) for which estimating an output probability with a large enough number of clicks is worst-case #P-hard, in direct analogy to the construction of the worst-case matrix AA above. Since this lies beyond the scope of the presented work, we leave this as future work.
@@
More specifically, we consider 256256-mode random unitary circuits drawn from local random circuit ensemble $\mathcal{H}_{\mathcal{A}}$ over the $q$-Kaleidoscope circuit architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ with the repetition number $q=1,2,3$, and global Haar random circuit ensemble. First, we sample 500 unitary circuits from each of the circuit ensembles described above. For each of the randomly chosen circuits, on input $N=4,8,12,16$ photons in the first NN modes over 256256 modes, we sample 500 outcomes using the Clifford-Clifford algorithm for Boson Sampling [clifford2018classical]. Then, we count the number of collision-free outcomes over the 500 sampled outcomes and compute the average ratio of collision-free outcomes for each NN and circuit ensemble, where the average is taken over the 500500 sampled unitary circuits for each circuit ensemble.

### 06 Mar 2026 v4

@@ Appendix E About Cayley transform
where we denote $\bm{t}^{*}$ as an input configuration vector permuted by $\bm{P}^{-1}$ from the fixed input configuration $\bm{t}$ in Eq. (1), and $S_{N}$ is $N$-mode permutation group. One can easily check that the common denominator for $\prod_{j=1}^{N}\left[V(\theta)_{\bm{s},\bm{t}^{*}}\right]_{\sigma_{j},j}$ is $\prod_{i=1}^{m}q_{i}(\theta)]^{N}$, and it does not change with permutation $\sigma$. Let us define $Q(\theta)=[\prod_{i=1}^{m}|q_{i}(\theta)|^{2}]^{N}$, which is a degree $4mN$ polynomial in $\theta$. Then, $Q(\theta)$ serves as the common denominator for the output probability. Hence, the output probability can be represented as $p_{\bm{s}}(V(\theta))=\frac{P(\theta)}{Q(\theta)}$, with $P(\theta)$ also a degree $4mN$ polynomial in $\theta$. ∎
@@ Appendix G Proof of Lemma 3
Let $\bar{p}_{\bm{s}}(C)$ be the output probability distribution from the approximate sampler $\mathcal{S}$ with the given linear optical circuit $C$. Also, let $\mathcal{C}_{M,N}$ be the set of collision-free outcomes of Boson Sampling, for mode number $M$ and photon number $N$. Then $\bar{p}_{\bm{s}}(C)$ satisfies
@@
in polynomial time in $N$ and $\alpha^{-1}$ is in $\rm{BPP}^{\rm{NP}^{\mathcal{S}}}$. Using $\operatorname*{\mathbb{E}}_{\bm{s}\sim\mathcal{G}_{M,N}}[\bar{p}_{\bm{s}}(C)]=\binom{M}{N}^{-1}\sum_{\bm{s}\in\mathcal{C}_{M,N}}\bar{p}_{\bm{s}}(C)\leq\binom{M}{N}^{-1}$,
@@ Appendix H On average-case hardness including collision outcomes
More specifically, let the mode number $M=\Theta(N^{\gamma})$ for any $\gamma\geq 1$. Let us define $\mathcal{Q}_{M,N}$ as the uniform distribution over all the possible outcomes of Boson Sampling with $M$ modes and $N$ photons, such that each outcome $\bm{s}\sim\mathcal{Q}_{M,N}$ is an $M$-dimensional output configuration vector $\bm{s}=(s_{1},\dots,s_{M})$ with the constraint $\sum_{i=1}^{M}s_{i}=N$. Also, we denote the number of $s_{i}=1$ in $\bm{s}$ as the number of singletons throughout this appendix. Here, let us assume that there exists $N_{0}=cN^{\lambda}$ for a constant $c$ and $0<\lambda<1$ such that it is likely that the number of singletons is more than $N_{0}$; we later provide the intuition to justify this assumption at the end of this appendix.
Hereafter, we denote each $(M_{0},N_{0})$ and $(M,N)$ as the system sizes (mode number & photon number) for the worst-case hardness (i.e., Theorem 3) and average-case hardness (we aim to prove), respectively. More specifically, given a worst-case circuit $C_{0}$ in $M_{0}=\Omega(N_{0})$ mode circuit architecture $\mathcal{B}\mathcal{B}^{*}$ and a fixed $M_{0}$-mode and $N_{0}$-photon collision-free outcome $\bm{s}_{0}$ depicted in Theorem 3, by our assumption, one can find a photon number $N$ that satisfies $N_{0}=cN^{\lambda}$ for a constant $c$ and $0<\lambda<1$, and mode number $M=\Theta(N^{\gamma})$ for $\gamma\geq 1$, such that most of the outcomes over $\mathcal{Q}_{M,N}$ have at least $N_{0}$ singletons. In the following, we show a possible approach for worst-to-average-case reduction for general outcomes, by describing how one can estimate $p_{\bm{s}_{0}}(C_{0})$ to within desired imprecision given oracle access to the well-estimated values $p_{\bm{s}}(U)$ for average-case outcomes $\bm{s}\sim\mathcal{Q}_{M,N}$ and $M$-mode average-case circuits over $U\sim\mathcal{H}_{\mathcal{A}}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$.
@@
We highlight the necessary ingredients that need to be addressed for the successful worst-to-average-case reduction by the above process. First, to make sure that we can also well-estimate $p_{\bm{s}_{1}}(U_{0})$ with high probability over $\bm{s}_{1}$ via the oracle access, one needs to justify that the distribution of $\bm{s}_{1}$ is sufficiently close to the distribution $\mathcal{Q}_{M/2,N-N_{0}}$. To help understand how the distribution of $\bm{s}_{1}$ can be close to $\mathcal{Q}_{M/2,N-N_{0}}$, consider the limiting case that $N_{0}=0$. Here, given that $M$ is much larger than $N$ (but still can be $M=\Theta(N)$), sampling $\bm{s}_{1}$ is equivalent to sampling from $\mathcal{Q}_{M,N}$ and post-selecting outcomes that have all the non-zero elements in the first $M/2$ mode, which yields the uniform distribution $\mathcal{Q}_{M/2,N}$. Therefore, given that $M$ is much larger than $N$ and $N_{0}$ is sufficiently small relative to $N$, we expect that the distribution of $\bm{s}_{1}$ would not largely deviate from the distribution $\mathcal{Q}_{M/2,N-N_{0}}$.
@@
We now present a toy model that could offer an insight that, for a large portion of outcomes of Boson Sampling, a sufficient number of singletons would exist. In the following, we present the classical “balls-into-bins" problem where we throw $N$ balls into $M$ bins uniformly at random, which is indeed the classical analogue of Boson Sampling. We show that, given that $M=\Theta(N^{\gamma})$ with $\gamma\geq 1$, it is highly unlikely that the number of bins getting exactly one ball is less than $cN^{\lambda}$ for a constant $c$ and $0<\lambda<1$.
Let us first denote the random variable $X_{i}$ as the number of balls in $i$-th bin (such that $\sum_{i=1}^{M}X_{i}=N$), when we throw $N$ balls into $M$ bins uniformly at random. Here, our goal is to verify that over the experiments, it is highly unlikely to observe the outcomes that the number of $X_{i}=1$ is smaller than $cN^{\lambda}$. Since all the variables $X_{i}$ are correlated due to the constraint $\sum_{i=1}^{M}X_{i}=N$, instead, we consider i.i.d. random variable $Y_{i}$ each drawn from Poisson distribution $P(\frac{N}{M})$, such that $\operatorname*{\mathbb{E}}[\sum_{i=1}^{M}Y_{i}]=N$ and $\Pr[\sum_{i=1}^{M}Y_{i}=N]=e^{-N}\frac{N^{N}}{N!}=\Theta(\frac{1}{\sqrt{N}})$. The motivation for using Poissonian random variables is that the distribution of $(Y_{1},\dots,Y_{M})$ conditioned on the sum $\sum_{i=1}^{M}Y_{i}=N$ is equivalent to the distribution of the balls-into-bins problem $(X_{1},\dots,X_{M})$ (see, e.g., Ref. [ross2014introduction]). Accordingly, if it is exponentially unlikely that the number of $Y_{i}=1$ is smaller than $cN^{\lambda}$ as we sample $(Y_{1},\dots,Y_{M})$, this indicates that it is still exponentially unlikely that the number of $X_{i}=1$ is smaller than $cN^{\lambda}$ as we sample $(X_{1},\dots,X_{M})$, given that the post-selection probability for $\sum_{i=1}^{M}Y_{i}=N$ is at most inverse polynomial (i.e., $\Theta(\frac{1}{\sqrt{N}})$).
@@
We conclude this appendix by introducing an alternative approach, based on the closely related result in [bouland2023complexity] that establishes a hardness argument for general Boson Sampling in the saturated regime. Note that our convention on singletons closely parallels the notion of “clicks" introduced in [bouland2023complexity], which refers to the number of modes that are occupied by a non-zero number of photons. Analogous to our analysis, their key observation is that, even in the saturated regime, most outcomes have a sufficiently large number of clicks. Based on this observation, they show that, given an $N$-photon outcome $\bm{s}$ with collisions that has a sufficient number of clicks $N_{0}$, there exists a poly-time constructible matrix $A\in\{0,1\}^{N_{0}\times N}$ such that computing $\text{Per}(A_{\bm{s}})$ is worst-case #P-hard, for $A_{\bm{s}}$ being a submatrix of $A$ by repeating rows according to $\bm{s}$. Then they establish a worst-to-average-case reduction by perturbing $A$ into their average-case matrix and performing polynomial interpolation (for details, see [bouland2023complexity]).
Accordingly, since we employ a conceptually similar strategy in the main text, establishing worst-case hardness for shallow-depth circuits with a large enough number of clicks is sufficient to obtain an analogous hardness result for shallow-depth Boson Sampling. This requires constructing a shallow-depth circuit (as done in the proof of Theorem 3) for which estimating an output probability with a large enough number of clicks is worst-case #P-hard, in direct analogy to the construction of the worst-case matrix $A$ above. Since this lies beyond the scope of the presented work, we leave this as future work.
@@
More specifically, we consider $256$-mode random unitary circuits drawn from local random circuit ensemble $\mathcal{H}_{\mathcal{A}}$ over the $q$-Kaleidoscope circuit architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ with the repetition number $q=1,2,3$, and global Haar random circuit ensemble. First, we sample 500 unitary circuits from each of the circuit ensembles described above. For each of the randomly chosen circuits, on input $N=4,8,12,16$ photons in the first $N$ modes over $256$ modes, we sample 500 outcomes using the Clifford-Clifford algorithm for Boson Sampling [clifford2018classical]. Then, we count the number of collision-free outcomes over the 500 sampled outcomes and compute the average ratio of collision-free outcomes for each $N$ and circuit ensemble, where the average is taken over the $500$ sampled unitary circuits for each circuit ensemble.